The dimensions of a certain computer monitor screen are such that its length is 2 inches more than its width. If the length is increased by 1 inch while the width remains the same, the area is increased by 16 square inches. What is the length and width of the screen
Answer by macston(5194)

.
W=width; L=length=W+2 A=Area
(L+1in)(W)=A+16 sq in
WL+Win=A+16sq in
WL=A Subtract from equation above
Win=16sq in Divide each side by 1 in
W=16 in ANSWER 1:The width is 16 inches.
L=W+2in=16in+2in=18 in: ANSWER 2: The length is 18 inches.
CHECK:
(L+1in)(W)=LW+16sq in
(18in+1in)(16in)=(18in)(16in)+16 sq in
(19in)(16in)=288 sq in+16 sq in
304 sq in=304 sq in